An economy has a labor endowment of 24 units. Good x has the production function x...

  1. An economy has a labor endowment of 24 units. Good x has the production function x = 10Lx and good y has the production function y = 2(Ly)0.5, where Lx and Ly are the labor units allocated to sector x and y, respectively. Find output x if a third of labor endowment goes to producing good x.
  1. 65
  2. 70
  3. 72
  4. 80
Answer #1

Answer

The correct answer is (d) 80.

Production function of x is given by :

x = 10Lx

where Lx = Labor units allocated to sector x

Here it is given that a third of labor endowment goes to producing good x and also it is given that Labor endowment = 24. Thus amount of labor goes to production of x = a third of 24 = 24/3 = 8

Hence Lx = 8

=> x = 10Lx = 10*8 = 80.

Thus,  output x if a third of labor endowment goes to producing good x = 80

Hence, the correct answer is (d) 80.